chloeharrellsk8:

The largest blimps ever built could carry 200,000 meters of hydrogen gas how many cubic decimeters is this?

luhivqqcherry:

is there options ?? .

Manny300303199:

Is the meters just meter or is it cubed

surjithayer:

1 meter=10 decimeter 1 cubic meter=10*10*10=1000 cubic decimeter 200,000 cubic meters=200,000*1000=200,000,000 cubic decimeters

toga:

To convert meters to decimeters, we multiply the value in meters by 10. Therefore, 200,000 meters is equal to 2,000,000 decimeters. To convert decimeters to cubic decimeters, we need to cube the value in decimeters. Therefore, 2,000,000 decimeters is equal to 8,000,000,000 cubic decimeters.